OpenStudy (twizttiez):

Use point slope form to write the equation of the line that passes through the given point and has the given slope. (1,2);m=-5

OpenStudy (twizttiez):

@ttb123456789

OpenStudy (anonymous):

yeah

OpenStudy (twizttiez):

I'm lost on this one

OpenStudy (anonymous):

y=-5x+7

OpenStudy (godlovesme):

y-y1= m(x-x1) y1= 2 x1=1 m= -5 can u solve the rest?

OpenStudy (godlovesme):

yea @ttb123456789 is right :D

OpenStudy (twizttiez):

Um sorry i asked the wrong question this is what i mean't to ask (0,0);m=3/4

OpenStudy (anonymous):

y=3/4

OpenStudy (anonymous):

it cannot be written in point slope

OpenStudy (twizttiez):

Ok 1 more

OpenStudy (anonymous):

yeah

OpenStudy (twizttiez):

(7,-8);M=1/7

OpenStudy (anonymous):

y=1/7x-9

OpenStudy (godlovesme):

y=1/7x-9 i guess :)

OpenStudy (twizttiez):

Thank you both ^_^

OpenStudy (godlovesme):

anytime...xD :D